On December 31, 2013, Ironman Steel issued $1,300,000, 8-year bonds for $1,340,000. The stated rate of interest was 7% and interest is paid annually on December 31.

Prepare the amortization table for Ironman Steel's bonds.

Period

Cash Payment (Credit)

Interest Expense (Debit)

Premium on Bonds Payable (Debit)

Premium on Bonds Payable Balance

Carrying Value
Answer #1

Amortization table

Period Cash payment (credit) Interest expense (debit) Premium on bonds payable (debit) Premium on bonds payable balance Carrying value
at issue 1340000
Dec 31,2014 1300000*7% = 91000 91000-5000 = 86000 40000/8 = 5000 35000 1335000
Dec 31,2015 91000 86000 5000 30000 1330000
Dec 31,2016 91000 86000 5000 25000 1325000
Dec 31,2017 91000 86000 5000 20000 1320000
Dec 31,2018 91000 86000 5000 15000 1315000
Dec 31,2019 91000 86000 5000 10000 1310000
Dec 31,2020 91000 86000 5000 5000 1305000
Dec 31,2021 91000 86000 5000 0 1300000
Total 728000 688000 40000